# From Grunts to Grammar: Emergent Language from Cooperative Foraging

### v1

@@ 4 Experimental Setting
4.1 Embodied Social Agents
This section provides an overview of how agents interact with their partners and the environment, and how they are trained. A formal description is also given in Appendix A. As shown in 2(b), each agent in our environment learns both to produce actions and to send discrete messages through a policy trained with PPO schulman2017proximal . We denote the overall policy of an agent as $\psi=(\pi,\phi)$, where $\pi$ is the action policy that selects an environment action, and $\phi$ is the communication policy that selects a discrete message $\mathbf{m}^{(t)}$ to send. We formulate communication as a sequence of discrete messages. At each time step $t$, an agent is only allowed to send one discrete message, where each message is represented by a single integer $\mathbf{m}^{(t)}\in\{0,1,2,3\}$. This integer indexes into a lookup table of 4 learnable embeddings (each of dimension 16) in the receiving agent. Each agent maintains its own message lookup table, which is not shared. In other words, upon receiving a message $\mathbf{m}^{(t)}$, the agent retrieves the corresponding message embedding from its own lookup table.
@@ 4.2 Training Regimes
We train each agent independently using standard single-agent PPO, with no shared parameters or centralized critic. Each agent treats its partner as part of the environment and learns solely from its own experience. Communication policies emerge through interaction and reward, without explicit supervision. This decentralized setup is intentional: it allows agents to specialize, diverge, and adapt to others, enabling us to study how communication protocols evolve under varying population sizes, agent heterogeneity, and connectivity patterns among agents. To investigate how communication strategies emerge and generalize in populations, we explore two training regimes that differ in how agents interact with other partners. These regimes allow us to study the effects of population diversity, message alignment, and exposure to self-produced messages on communication development.
Cross-Play Training (XP): We define the agent population size as $N_{\text{pop}}$. In each training episode, a pair of agents is randomly sampled from the population to play together.
Cross-Play-and-Self-Play Training (XP+SP): Humans can speak to themselves and understand their own language. Motivated by this aspect of human learning, in each episode, we randomly select either two distinct agents (Cross-Play, XP) or the same agent twice (Self-Play, SP). Specifically, for each episode, we randomly sample agents $i$ and $j$ such that either $i\neq j$ or $i=j$, where $i,j\in\{1,\dots,N_{\text{pop}}\}$.
4.3 Social Network Structure
We investigate two social network structures, as illustrated in 4(a) and 4(b): fully connected (FC) and ring-structured (Ring) networks. In the fully connected setup, each agent interacts with all other agents during training, representing a standard approach in multi-agent systems. In contrast, the ring-structured population limits each agent to interacting only with its immediate neighbors, forming a circular network. This structure is chosen because it provides an extreme scenario where agents have the minimal number of neighbors, yet still maintain connectivity between all agents through indirect paths. This setup enables us to examine how agents generalize to unseen partners and how communication protocols develop in a more constrained, sparsely connected environment.
4.4 Evaluation Metrics on Language Properties
To evaluate linguistic properties in emergent communication, we use two standard metrics, Topographic Similarity brighton2006understanding ; lazaridou2018emergence and Language Similarity kim2021emergent ; rita2022role ; michelrevisiting . We also introduce a new metric, Interchangeability below. Formal definitions of these metrics are provided in Appendix D.
Topographic Similarity (topsim) measures the structural alignment between message space and semantic space. It is computed as the Spearman correlation between pairwise distances in the message space and the corresponding distances in the semantic space, which reflects the agent’s observations and context. A higher topsim indicates a more compositional and consistent mapping between semantic meanings and messages. We define the semantic space using ground-truth attributes of the environment, specifically the scores and positions of items in the ScoreG game of FG. To implement the topographic similarity metric, we use the EGG toolkit kharitonov2019egg .
Language Similarity (LS) measures how similarly two agents communicate in the same situation. It compares the sequences of discrete messages each agent produces and computes the average agreement between them. A higher score means the agents tend to use the same messages in similar contexts, indicating stronger convergence in their communication strategies.
Interchangeability (IC) is a metric we propose to measure whether an agent can understand the language it produces. We define Self-SR as the success rate when an agent plays with a copy of itself, and Cross-SR as the success rate when paired with a different agent. The overall success rate across both settings is denoted as SR. IC is the ratio of Cross-SR to Self-SR. A high IC indicates that agents can better generalize and interpret their own language.
4.5 Implementation Details
@@
All training can be performed on a single GeForce RTX 4080 (See Appendix Figure S4). We use the ADAM optimizer kingma2014adam with an initial learning rate of $0.00025$, which linearly decays over time. We report the hyperparameters of the architecture and training algorithm in Appendix G.
@@
5 Results
Table 1: Cross-play training can cause non-interchangeable languages. Game performance and language similarity (LS; Sec. 4.4) comparison across different training regimes. $N_{\text{pop}}$ is the population size. Cross-SR and Self-SR are the mean ($\pm$ standard deviation) success rates when agents play with others or with copies of themselves, respectively.
@@ 5 Results
(a) topsim across population sizes
(b) IC across population sizes
(c) LS across population sizes
Figure 3: The effect of population training with fully-connected social networks on language properties. (a) Topographic Similarity (topsim), (b) Interchangeability (IC), and (c) Language Similarity (LS) as a function of population sizes are shown. See Sec. 4.4 for metrics details.
@@
We investigate the emergence of linguistic properties in the FG framework and present our findings in this section. Each experimental setting is defined by four factors: the game type, agent population size, social structure, and training regime. These factors influence the emergence of different language properties, making exhaustive analysis infeasible. Instead, we report results from a representative subset of settings. For clarity, we adopt the naming convention for each experimental setting: [game]-[population size]-[social structure]-[training regime]. For example, ScoreG-P15-Ring-XP+SP refers to a setting where $N_{\text{pop}}=15$ agents in a ring-structured network are trained with XP+SP regime on the ScoreG game. This convention is used throughout our analysis.
@@
Increasing population size and self-play training support interchangeable language [ScoreG-FC]. As shown in Table 1, XP agents trained exclusively with each other fail to understand their own messages when paired with a copy of themselves during test time as indicated by a drop of 6% in Self-SR. Surprisingly, when agents are trained with $N_{\text{pop}}=3$, they maintain high Self-SR even without explicit self-play, suggesting that population training encourages the emergence of interchangeable language. These agents also develop more consistent language, as indicated by a higher LS of 0.53 compared to the XP alone. Furthermore, XP+SP agents achieve high Self-SR due to their direct exposure to SP. Interestingly, they also exhibit greater language similarity across agents, with an LS of 0.59, outperforming XP agents even in population settings. We also provide a theoretical analysis to explain why training with more partners and training with SP can lead to a shared language in Appendix B.
@@
Population size affects compositionality [ScoreG-FC]. As shown in 3(a), population size has an effect on topsim. For XP agents, topsim increases from population size 2 and saturates from size 6. In contrast, XP+SP agents show a more consistent upward trend in topsim as population size increases, plateauing around size 12–15. Interestingly, XP+SP agents trained in a population of size 2 also achieve the highest topsim across all population sizes. However, we caution against the interpretation of effective language from such a small population because this language could be overfitting to a fixed partner and may not generalize well to more diverse partners. Linguistic and cognitive studies suggest that larger communities tend to produce more compositional languages reali2018simpler ; raviv2019larger . While population training promotes consistency, we observe that compositionality does not increase monotonically, aligned with previous works kim2021emergent ; rita2022role ; michelrevisiting .
@@
Population size affects language similarity and interchangeability [ScoreG-FC]. 3(b) shows that IC quickly saturates to nearly 1.0 for XP agents once the population size exceeds 2. This indicates that agents can understand their own messages. For XP+SP agents, IC reaches the maximum level across $N_{\text{pop}}\geq 2$. This indicates that SP encourages language interchangeability. As shown in 3(c), for XP agents, LS increases sharply from size 2 to 3, then gradually plateaus, indicating that training with multiple partners encourages the emergence of a shared protocol. XP+SP agents, in contrast, show consistently high LS across all population sizes. This suggests that self-play helps promote consistent, shared communication protocols within a population. Together, these results suggest that both self-play and population diversity contribute to the emergence of interchangeable and shared communication. Unlike our findings, prior work kim2021emergent reported that larger population sizes reduce language similarity, as listeners adapt to multiple speaker-specific languages. However, this discrepancy may stem from their unidirectional communication setup and centralized agent training, resulting in different optimal solutions than ours.
@@
Self-play supports interchangeable and generalizable language in ring-structured populations [ScoreG-P15-Ring]. We investigate how language propagates through a structured population by training agents in a ring topology which is the simplest non-fully connected social network kim2021emergent ; michelrevisiting . Specifically, we set the population size to $N_{\text{pop}}=15$ and train each agent to interact only with its two neighbors. This setup enables us to examine how culture gets transmitted across indirect connections within the population at the test time. In Figure 4, for XP+SP agents, both LS and SR gradually decline as the distance between agents increases. This indicates that the learned language generalizes to indirectly connected agents, though with reduced SR. In contrast, for agents with XP alone, language becomes non-interchangeable across the population. For example, an agent may produce messages more similar to those of a non-neighbor than to its direct partner at the test time. This asymmetry implies that agents learn to understand their neighbors but fail to produce messages aligned with the language they understand (2(c)). Together, these results suggest that self-play not only improves within-agent consistency but also supports more stable and transmissible communication protocols across structured populations.
@@
Spatial and temporal displacement emerge in agent communication [P3-FC-XP]. Humans can refer to past or future events—for example, I saw an apple yesterday or I will collect an apple tomorrow. They can also refer to spatially distant objects, such as an apple in the forest. This ability to refer to things beyond the immediate here and now is known as displacement in natural language. We investigate whether displacement can emerge in artificial foraging agents using the TemporalG game (section 3). XP agents with $N_{\text{pop}}=3$ achieve strong performance in this setting, with a Cross-SR of $0.975$ and a Self-SR of $0.962$ (Table S1). Next, we decode spatial and temporal displacement information from the messages. The items can take on one of 6 possible spawn times, 4 vertical positions (excluding the grid’s center), and 5 possible horizontal positions. The decoding accuracy for the spawn time and position of items is above chance (5(b)), suggesting that messages function as temporal adverbials and spatial references. We also decode items’ scores and positions from XP agents ($N_{\text{pop}}=3$) in the ScoreG game (5(a)). The items can take on one of 10 possible score ranges, 2 possible vertical positions (top and bottom), and 5 possible horizontal positions. The results align with those observed in the TemporalG game. Moreover, compared to Integer Msg, decoding with chains of message embeddings yields higher accuracy, suggesting that these embeddings encode more meaningful and linearly separable spatial and temporal features.
(a) Fully-Connected
(b) Ring-Structured
(c) LS across circular distance
(d) SR across circular distance
Figure 4: Comparison of population training ($N_{\text{pop}}=15$) with self-play (XP+SP) and without self-play (XP) under ring-structured social networks: (a,b) are social network structures described in Sec. 4.3 (b,c) We plot LS and SR (See Sec. 4.4) as functions of circular distances between two agents in a ring-structured network. A distance of 1 indicates that two agents were co-trained, while a distance greater than 1 indicates that the agents were never paired during training.
@@
Implicit communication can emerge when explicit message communication is disabled [ScoreG-P3-FC-XP]. We study whether agents can convey information through their sequence of actions in an episode when the explicit communication channel is disabled. We conduct an ablation on two variables: partner visibility and the presence of explicit verbal communication. When a partner is invisible, their location appears as an unoccupied cell in the observing agent’s receptive field. We train XP agents in Inv-NoCom and Vis-NoCom, where agents either have or lack partner visibility, but no explicit message communication is allowed. If an agent observes an extreme score (e.g., 222 or 22), it can confidently infer whether the item is a goal or not. To isolate this confounding factor, we evaluate agents only under the condition that both partners observe high scores. Inv-Com agents (XP agents trained in our default FG environments) achieve SR of $85\%$. Without communication, Inv-NoCom agents perform below chance ($40\%$), likely because they fail to coordinate goal pickup without seeing each other. Interestingly, Vis-NoCom agents achieve an SR of 60%, outperforming Inv-NoCom. This indicates that both Inv-Com and Vis-NoCom agents can glean information from their partners—either via explicit messages or by observing partners’ actions—consistent with findings in grupen2022multi ; wang2025scout . We also evaluate the average length of successful test episodes. Both Inv-Com and Vis-NoCom agents take around 6 steps, while Inv-NoCom agents take about 5 steps. This indicates that agents with more accurate communication require extra steps to send meaningful signals, increasing their communication bandwidth and hence, longer episode length.
@@
(a) Success Rate
(b) Avg Ep Length
Figure 6: Implicit communication in the ScoreG games. Inv/Vis denote partner visibility; Com/NoCom indicate presence or absence of verbal communication. Item scores are from $[160,162,\dots,240]$. We report average successful episode length; error bars show standard deviations across seeds and agents.

### v2

@@ 4 Experimental Setting
Embodied Social Agents. As shown in 2(b), each agent in our environment learns both to produce actions and to send and receive discrete messages through a policy trained with PPO schulman2017proximal (Sec. A.1). For an experiment on unidirectional communication, see Sec. D.2. We denote the overall policy of an agent as $\psi=(\pi,\phi)$, where $\pi$ is the action policy that selects an environment action, and $\phi$ is the communication policy that selects a discrete message $\mathbf{m}^{(t)}$ to send. We formulate communication as a sequence of discrete messages. At each time step $t$, an agent is only allowed to send one discrete message, where each message is represented by a single integer $\mathbf{m}^{(t)}\in\{0,1,2,3\}$. This integer indexes into a lookup table of 4 learnable embeddings (each of dimension 16) in the receiving agent. Each agent maintains its own message lookup table, which is not shared. In other words, upon receiving a message $\mathbf{m}^{(t)}$, the agent retrieves the corresponding message embedding from its own lookup table. We also report results for vocabulary sizes other than the default of 4 in Table S5.
@@
Training Regimes. We train each agent independently using standard single-agent PPO, with no shared parameters or centralized critic. Each agent treats its partner as part of the environment and learns solely from its own experience. Communication policies emerge through interaction and reward, without explicit supervision. This decentralized setup is intentional: it allows agents to specialize, diverge, and adapt to others, enabling us to study how communication protocols evolve under varying population sizes, agent heterogeneity, and connectivity patterns among agents. To investigate how communication strategies emerge and generalize in populations, we explore two training regimes that differ in how agents interact with other partners. These regimes allow us to study the effects of population diversity, message alignment, and exposure to self-produced messages on communication development. Cross-Play Training (XP): We define the agent population size as $N_{\text{pop}}$. In each training episode, a pair of agents is randomly sampled from the population to play together. Cross-Play-and-Self-Play Training (XP+SP): Motivated by this aspect of inner monologue, humans being able to speak to themselves, in each episode, we randomly select either two distinct agents (Cross-Play, XP) or the same agent twice (Self-Play, SP). Specifically, for each episode, we randomly sample agents $i$ and $j$ such that either $i\neq j$ or $i=j$, where $i,j\in\{1,\dots,N_{\text{pop}}\}$.
Social Network Structure. We consider several social network structures (Figure S3): Fully Connected (FC), where every agent interacts with all others; Ring, where each agent interacts only with immediate neighbors; Ring with Cliques (Clq), which adds short-range clusters to the ring; Watts–Strogatz (WS), which rewires a fraction of ring edges to create small-world shortcuts; and Ring with Long-Range Connections (LRC), which augments the ring with distant edges. FC represents the dense baseline. Ring is the sparsest structure while still connected, allowing us to test how communication propagates along long paths. Adding cliques or long-range links (Clq, WS, LRC) reduces social distance and is expected to strengthen language transmission compared to Ring ( Appendix C).
Evaluation Metrics on Language Properties. To evaluate linguistic properties in emergent communication, we use two standard metrics and a new metric below. Formal definitions of these metrics are provided in Sec. A.3. Briefly, Topographic Similarity (topsim) brighton2006understanding ; lazaridou2018emergence measures the structural alignment between message space and semantic space. It is computed as the Spearman correlation between pairwise distances in the message space and the corresponding distances in the semantic space, which reflects the agent’s observations and context. A higher topsim indicates a more compositional and consistent mapping between semantic meanings and messages. Language Similarity (LS) kim2021emergent ; rita2022role ; michelrevisiting measures how similarly two agents communicate in the same situation. It compares the sequences of discrete messages each agent produces and computes the average agreement between them. Interchangeability (IC) measures whether an agent can understand the language it produces. We define Self-SR as the success rate when an agent plays with a copy of itself, and Cross-SR as the success rate when paired with a different agent. The overall success rate is denoted as SR. IC is the ratio of Self-SR to Cross-SR. A high IC indicates that agents can better generalize and interpret their own language.
@@
Implementation Details. All training can be performed on a single GeForce RTX 4090 (See Appendix Figure S1 and Figure S2). We use the ADAM optimizer kingma2014adam with an initial learning rate of $0.00025$, which linearly decays over time. We report the hyperparameters of the architecture and training algorithm in Sec. A.2 (Table S1 and Table S2).
@@
5 Result
Table 1: Cross-play training can cause non-interchangeable languages. Game performance and language similarity (LS; Sec. A.3) comparison across different training regimes. $N_{\text{pop}}$ is the population size. Cross-SR and Self-SR are the mean ($\pm$ standard deviation) success rates when agents play with others or with copies of themselves, respectively.
@@ 5 Result
We investigate the emergence of linguistic properties in the FG framework according to previously stated research questions (Q1-Q5) and present our findings in this section. Each experimental setting is defined by four factors: the game type, agent population size, social structure, and training regime. These factors influence the emergence of different language properties, making exhaustive analysis infeasible. Instead, we report results from a representative subset of settings. For clarity, we adopt the naming convention for each experimental setting: [game]-[population size]-[social structure]-[training regime]. For example, ScoreG-P15-Ring-XP+SP refers to a setting where $N_{\text{pop}}=15$ agents in a ring-structured network are trained with XP+SP regime on the ScoreG game. This convention is used throughout our analysis.
@@
Increasing population size and self-play training support interchangeable language [ScoreG-FC]. Since decentralized agents do not share neural network parameters, it is natural to wonder whether they develop a common language at all. A curious possibility is that one agent might produce L1 while only understanding L2, with the other doing the opposite. Moreover, it is possible that an agent trained with its partner does not understand the language it produces. As shown in Table 1, XP agents trained exclusively with each other fail to understand their own messages when paired with a copy of themselves during test time as indicated by a drop of 6% in Self-SR. Surprisingly, when agents are trained with $N_{\text{pop}}=3$, they maintain high Self-SR even without explicit self-play, suggesting that population training encourages the emergence of interchangeable language. These agents also develop more consistent language, as indicated by a higher LS of 0.53 compared to the XP alone. Furthermore, XP+SP agents achieve high Self-SR due to their direct exposure to SP. Interestingly, they also exhibit greater LS across agents, with an LS of 0.59, outperforming XP agents even in population settings. Agents with high LS produce similar messages, while agents with low LS produce distinct ones in the embedding space (Figure S4). Finally, we provide a theoretical analysis showing why training with more partners and with SP leads to more common language in Appendix B.
@@
Population size affects compositionality [ScoreG-FC]. Linguistic and cognitive studies suggest that larger communities are more likely to develop compositional languages reali2018simpler ; raviv2019larger . Motivated by this, we investigate how compositionality varies with increasing population size. As shown in 3(a), population size has an effect on topsim. For XP agents, topsim increases from population size 2 and saturates from size 6. In contrast, XP+SP agents show a more consistent upward trend in topsim as population size increases, plateauing around size 12–15. Interestingly, XP+SP agents trained in a population of size 2 also achieve the highest topsim across all population sizes. However, we caution against the interpretation of effective language from such a small population because this language could be overfitting to a fixed partner and may not generalize well to more diverse partners. While population training promotes consistency, we observe that compositionality does not increase monotonically, aligned with results reported in previous studies kim2021emergent ; rita2022role ; michelrevisiting .
(a) topsim across $N_{\text{pop}}$
(b) IC across $N_{\text{pop}}$
(c) LS across $N_{\text{pop}}$
Figure 3: The effect of population training with fully-connected social networks on language properties. (a) Topographic Similarity (topsim), (b) Interchangeability (IC), and (c) Language Similarity (LS) as a function of population sizes are shown. See Sec. A.3 for metrics details.
@@
Population size affects language similarity and interchangeability [ScoreG-FC]. We observed that a population of three agents can develop a shared, interchangeable language. We now ask whether this effect persists as the population size increases. 3(b) shows that IC quickly saturates to nearly 1.0 for XP agents once the population size exceeds 2. This indicates that agents can understand their own messages. For XP+SP agents, IC reaches the maximum level across $N_{\text{pop}}\geq 2$. As shown in 3(c), for XP agents, LS increases sharply from size 2 to 3, then gradually plateaus, indicating that training with multiple partners encourages the emergence of a shared protocol. XP+SP agents, in contrast, show consistently high LS across all population sizes. This suggests that self-play helps promote consistent, shared communication protocols within a population. Together, these results suggest that both self-play and population diversity contribute to the emergence of interchangeable and shared communication. Unlike our findings, prior work kim2021emergent reported that larger population sizes reduce language similarity, as listeners adapt to multiple speaker-specific languages. This discrepancy may stem from their unidirectional communication and centralized training, resulting in different optimal solutions.
@@
Self-play enhances cultural transmission in weakly-connected networks [ScoreG-P15]. Human language evolves in sparsely connected societies, giving rise to a shared language alongside dialects and foreign languages. To investigate how language propagates in such settings, we train agents within weakly connected Ring and Clq social networks. The population size is fixed at $N_{\text{pop}}=15$, with each agent interacting only with a few neighbors. This setup allows us to examine how culture is transmitted across indirect connections at test time. As shown in Figure 4, both LS and SR generally decline with increasing distance between agents. This indicates that language partially propagates through non-co-trained agents: nearby neighbors develop more similar languages than distant ones. Importantly, the learned language also generalizes to indirectly connected agents, albeit with reduced SR. An exception arises for XP agents under the Ring network, which display a distinctive zig-zag pattern in 4(a) and 4(b). Here, an agent produces messages more similar to those of a non-neighbor than to its direct partner (5(c)). While agents successfully learn to interpret their neighbors, they fail to produce messages aligned with the language they understand (5(c) and 6(c)), highlighting a breakdown in interchangeability. Furthermore, as shown in 4(a) and 4(c), self-play training yields consistently higher LS across all social distances compared to cross-play training. Together, these findings suggest that self-play not only enhances within-agent consistency but also supports more stable and transmissible communication protocols in structured populations. Finally, we find that adding more connections to Ring can greatly enhance cultural transmission (Figure S7).
@@
Spatial and temporal displacement emerge in agent communication [P3-FC-XP]. Humans can refer to past or future events, for example, I saw an apple yesterday or I will collect an apple tomorrow. They can also refer to spatially distant objects, such as an apple in the forest. This ability to refer to things beyond the immediate here and now is known as displacement in natural language. We investigate whether displacement can emerge in artificial foraging agents using the TemporalG game. XP agents with $N_{\text{pop}}=3$ achieve strong performance in this setting, with a Cross-SR of $0.975$ and a Self-SR of $0.962$ (Table S11). Since the communication range is limited, the agents adopt a rendezvous strategy, meeting near the center of the grid map to establish communication (Figure S9). Next, we decode spatial and temporal displacement information from the messages. The items can take on one of 6 possible spawn times, 4 vertical positions (excluding the grid’s center), and 5 possible horizontal positions. The decoding accuracy for the spawn time and position of items is above chance (5(b)), suggesting that messages function as temporal adverbials and spatial references. We also decode items’ scores and positions from XP agents ($N_{\text{pop}}=3$) in the ScoreG game (5(a)). The items can take on one of 10 possible score ranges, 2 possible vertical positions (top and bottom), and 5 possible horizontal positions. The results align with those observed in the TemporalG game. Moreover, compared to Integer Msg, decoding with chains of message embeddings yields higher accuracy, suggesting that these embeddings encode more meaningful and linearly separable spatial and temporal features. Interestingly, when agents collect targets remotely without movement, their messages stop conveying spatial information (Figure S8), highlighting the role of embodiment in shaping language.
(a) LS
(b) SR
(c) LS
(d) SR
Figure 4: Self-Play (SP) promotes cultural transmission across social networks ($N_{\text{pop}}=15$): We plot LS and SR (see Sec. A.3) as a function of the shortest-path distance between two agents. (a,b) ring-structured network (Ring). (c,d) ring-structured with clique network (Clq).
@@
Implicit communication can emerge when explicit message communication is disabled [ScoreG-P3-FC-XP]. We study whether agents can convey information through their sequence of actions in an episode when the explicit communication channel is disabled. We conduct an ablation on two variables: partner visibility and the presence of explicit verbal communication. When a partner is invisible, their location appears as an unoccupied cell in the observing agent’s receptive field. We train XP agents in Inv-NoCom and Vis-NoCom, where agents either have or lack partner visibility, but no explicit message communication is allowed. If an agent observes an extreme score (e.g., 222 or 22), it can confidently infer whether the item is a goal or not. To isolate this confounding factor, we evaluate agents only under the condition that both partners observe high scores. As shown in Figure 6, Inv-Com agents (XP agents trained in our default FG environments) achieve SR of $85\%$. Without communication, Inv-NoCom agents perform below chance ($40\%$), likely because they fail to coordinate goal pickup without seeing each other. See more failure analysis in Sec. D.6. Interestingly, Vis-NoCom agents achieve an SR of 60%, outperforming Inv-NoCom. This indicates that both Inv-Com and Vis-NoCom agents can glean information from their partners—either via explicit messages or by observing partners’ actions. We also evaluate the average length of successful test episodes. Both Inv-Com and Vis-NoCom agents take around 6 steps, while Inv-NoCom agents take about 5 steps. This indicates that agents with more accurate communication require extra steps to send meaningful signals, increasing their bandwidth and hence, longer episode length. Another possible strategy for implicit communication is bumping into each other; however, our analysis in Sec. D.7 shows this is not the case. The similar analysis above on implicit communication in ScoreG is also applicable in TemporalG (Sec. E.1).
@@
(c) Success Rate
(d) Avg Ep Length
Figure 6: Implicit communication in the ScoreG games. Inv/Vis denote partner visibility; Com/NoCom indicate presence or absence of verbal communication. Item scores are from $160,162,\dots,240]$. We report average successful episode length; error bars show standard deviations across seeds and agents.